OpenStudy (anonymous):

Let R={0, 1, 2, 3} be the range of h(x) = x - 7. The domain of h inverse is A.{-7, -6, -5, -4} B.{7, 8, 9, 10} C.{0, 1, 2, 3}

OpenStudy (anonymous):

b

OpenStudy (anonymous):

get the inverse to check it out

OpenStudy (anonymous):

h^(-1) = x+7 0+7 = 7 ... 1 +7 = 8 ... etc...

OpenStudy (anonymous):

oops... * h^(-1) (x)